# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 2.4.2. Fusion estimates
We now turn to the fusion estimates, which are key to the factorisation of the residues as a product of primary field and Dotsenko-Fateev integral. For this, we need the radial decomposition of the free field: we write
$X(e^{-t+i\theta})=B_{t}+\varphi_{t}(e^{i\theta}),$
where $B_{t}=\int_{0}^{2\pi}X(e^{-t+i\theta})\frac{\mathrm{d}\theta}{2\pi}$ is a standard Brownian motion independent of $\varphi_{t}$. The field $e^{-t+i\theta}\mapsto\varphi_{t}(e^{i\theta})$ is $\log$-correlated in $\mathbb{D}$, and we define (after regularisation)
$Z_{t}:=\int_{\mathbb{A}_{t}}e^{\gamma\varphi_{t}(e^{i\theta})-\frac{\gamma^{2}}{2}\mathbb{E}[\varphi_{t}(e^{i\theta})^{2}]}\mathrm{d}t\mathrm{d}\theta$
the total GMC mass of the annulus $\mathbb{A}_{t}$ viewed as the cylinder $(0,t)\times\mathbb{S}^{1}$. Almost surely, the process $\mathbb{R}_{+}\ni t\mapsto Z_{t}$ is strictly increasing, so we can define its differential $\mathrm{d}Z_{t}$, which is a (random) measure on $\mathbb{R}_{+}$.
[…]
@@
Let $r\in\mathbb{N}^{*}$ and $\alpha\in\mathbb{R}$. The following estimates hold in $\mathcal{C}^{\prime}$.
@@
    Suppose $\alpha+r\gamma\;\geqslant\;Q$. Then,
@@
    Suppose $\gamma<\sqrt{2}$ and $\alpha+r\gamma<Q$. Then,
    $\Psi_{\alpha}(\boldsymbol{w})=\prod_{j=1}^{r}|w_{j}|^{-\gamma\alpha}\prod_{1\;\leqslant\;k<l\;\leqslant\;r}|w_{k}-w_{l}|^{-\gamma^{2}}\left(\Psi_{\alpha+r\gamma}+O(\max_{j}|w_{j}|^{\xi})\right),$
    for some $\xi>0$.
@@
By permutation symmetry, we can assume $|w_{1}|\;\geqslant\;|w_{2}|...\;\geqslant\;|w_{r}|$. Let $F\in\mathcal{C}$. By the Girsanov transform, we can write
$\Psi_{\alpha}(\boldsymbol{w})$ $=e^{(\alpha+r\gamma-Q)c}\prod_{j=1}^{r}|w_{j}|^{-\gamma\alpha}\prod_{1\;\leqslant\;k<l\;\leqslant\;r}|w_{k}-w_{l}|^{-\gamma^{2}}$
×$\qquad\times\mathbb{E}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}F\left(\varphi+\gamma\sum_{j=1}^{r}G(w_{j},\cdot\,)\right)\right]$
Our goal will be to bound the expecation in the last formula. We remark that the shifted function $F(\varphi+\gamma\sum_{j=1}^{r}G(w_{j},\cdot\,))$ converges to $F$ in $\mathcal{C}$ as $\boldsymbol{w}\to 0$, so we will omit this term. in our analysis.
@@
The circle average process of the field $X(z)-\gamma\sum_{j=1}^{r}\log|z-w_{j}|$ is $B_{t}+\gamma\sum_{j=1}^{r}t\wedge t_{j}$, where $t_{j}:=\log\frac{1}{|w_{j}|}$. By the Girsanov transform, we then have
$\mathbb{E}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}\right]$
$\quad=\mathbb{E}\left[e^{(\alpha+r\gamma-Q)B_{t_{1}}-\frac{1}{2}(\alpha+r\gamma-Q)^{2}}e^{-\mu e^{\gamma c}\int_{0}^{\infty}e^{\gamma B_{t}-u_{t}}\mathrm{d}Z_{t}}\right]$
$\quad\;\leqslant\;\frac{|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}}{(2\pi\log\frac{1}{|w_{1}|})^{1/2}}\mathbb{E}\left[\int_{\mathbb{R}}e^{(\alpha+r\gamma-Q)x}e^{-\mu e^{\gamma(c+x)}\int_{t_{1}}^{\infty}e^{\gamma(\widetilde{B}_{t-t_{1}}-u_{t})}\mathrm{d}Z_{t}}\mathrm{d}x\right]$
$\quad=\frac{1}{\gamma}\Gamma\left(\frac{1}{\gamma}(\alpha+r\gamma-Q)\right)\frac{|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}}{(2\pi\log\frac{1}{|w_{1}|})^{1/2}}\mathbb{E}\left[\left(\mu e^{\gamma c}\int_{t_{1}}^{\infty}e^{\gamma(\widetilde{B}_{t-t_{1}}-u_{t})}\mathrm{d}Z_{t}\right)^{-\frac{1}{\gamma}(\alpha+r\gamma-Q)}\right]$
$\quad=O(|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}).$
@@
We introduce the notation
$I_{D}(\boldsymbol{w})=\int_{D}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}},$
for every open set $D\subset\mathbb{D}$. It admits positive moments of order $p>0$ for all $p<\min\{\frac{4}{\gamma^{2}},\frac{2}{\gamma}(Q-\alpha-r\gamma)\}$ [KRV20, Equation (2.14)]. We will assume $p<1$ so that the condition is just $p<\frac{2}{\gamma}(Q-\alpha-r\gamma)$, since $\frac{4}{\gamma^{2}}>1$.
@@
Our goal is to control $\mathbb{E}[e^{-\mu I_{\mathbb{D}}(\boldsymbol{w})}-e^{-\mu I_{\mathbb{D}}(0)}]$. First, we show that we can remove the contribution of the disc of radius $|w_{1}|^{1-\eta}$ around 0, for any $\eta\in(0,1)$. By Hölder regularity of the exponential, we have
$\mathbb{E}\left[e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})}-e^{-\mu I_{\mathbb{D}}(\boldsymbol{w})}\right]$ $=\mathbb{E}\left[e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})}\left(1-e^{-\mu I_{|w_{1}|^{1-\eta}\mathbb{D}}}\right)\right]$
$\;\leqslant\;\mathbb{E}\left[1-e^{-\mu I_{|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})}\right]$
$\;\leqslant\;C\mathbb{E}[(I_{|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w}))^{p}]=O(|w_{1}|^{(1-\eta)\xi(p)}),$
for some $C>0$ and a multifractal exponent $\xi(p)>0$. Similarly, we get $\mathbb{E}[e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)}-e^{-\mu I_{\mathbb{D}}(0)}]=O(|w_{1}|^{(1-\eta)\xi(p)})$.
[…]
$\mathbb{E}\left[\left(I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})-I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)\right)e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)}\right]$ (2.6)
$\quad=\int_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}\left(\prod_{j=1}^{r}|w-w_{j}|^{-\gamma^{2}}|w|^{-\gamma\alpha}-|w|^{-\gamma(\alpha+r\gamma)}\right)\mathbb{E}\left[e^{-\mu\int_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}|z-w|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma(\alpha+r\gamma)}}}\right]|\mathrm{d}w|^{2}.$
From here, we need to distinguish two cases.
Case $\alpha+r\gamma<\frac{2}{\gamma}$.
The singularity $|w|^{-\gamma(\alpha+r\gamma)}$ is integrable with respect to Lebesgue measure $|\mathrm{d}w|^{2}$. We can also bound the expectation in (2.6) by 1. Therefore, (2.6) is bounded by a constant times
$\int_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}\Big{|}|w|^{-\gamma(\alpha+r\gamma)}-\prod_{j=1}^{r}|w-w_{j}|^{-\gamma^{2}}|w|^{-\gamma\alpha}\Big{|}|\mathrm{d}w|^{2}$ $=\int_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}\Big{|}1-\prod_{j=1}^{r}\left(1-\frac{|w_{j}|}{|w|}\right)^{-\gamma^{2}}\Big{|}\frac{|\mathrm{d}w|^{2}}{|w|^{\gamma(\alpha+r\gamma)}}$
$\quad\;\leqslant\;C|w_{1}|^{\eta}\int_{\mathbb{D}}\frac{|\mathrm{d}w|^{2}}{|w|^{\gamma(\alpha+r\gamma)}}=O(|w_{1}|^{\eta}).$
where we have just used $\frac{|w_{j}|}{|w|}\;\leqslant\;|w_{1}|^{\eta}$ to bound the integrand by $C|w_{1}|^{\eta}$ for some $C>0$.
Case $\alpha+r\gamma\;\geqslant\;\frac{2}{\gamma}$.
Note that $\alpha+(r+1)\gamma\;\geqslant\;\frac{2}{\gamma}+\gamma>Q$, so the singularity $|z|^{-\gamma(\alpha+(r+1)\gamma)}$ is not integrable with respect to $M_{\gamma}$. Using the same method as in the proof of Item 1., the expectation in (2.6) is $O(|w_{1}|^{\frac{1}{2}(\alpha+(r+1)\gamma-Q)^{2}}$. Hence (2.6) is bounded by a constant times
$\int_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}\left(|w|^{-\gamma(\alpha+r\gamma)}-\prod_{j=1}^{r}|w-w_{j}|^{-\gamma^{2}}|w|^{-\gamma\alpha}\right)|w|^{\frac{1}{2}(Q-\alpha-(r+1)\gamma)^{2}}|\mathrm{d}w|^{2}$
$\quad\;\leqslant\;C|w_{1}|^{\eta}\int_{\mathbb{D}}|w|^{-\gamma(\alpha+r\gamma)+\frac{1}{2}(\alpha+(r+1)\gamma-Q)^{2}}|\mathrm{d}w|^{2}$
Observe that $-\gamma(\alpha+r\gamma)+\frac{1}{2}(\alpha+(r+1)\gamma-Q)^{2}=\frac{1}{2}(\alpha+r\gamma-Q)^{2}-2>-2$, so that the last integral is indeed finite. ∎
Remark 1.
Although we won’t be needing this, the condition $\gamma<\sqrt{2}$ can be removed in item 2. Indeed, the only time this condition is required is for the last two displays of the proof to be finite (integrability of $|w-w_{j}|^{-\gamma^{2}}$). However, in the case $\gamma\;\geqslant\;\sqrt{2}$, we have the freezing estimate $O(|w-w_{j}|^{-\gamma^{2}+\frac{1}{2}(Q-2\gamma)^{2}})$ as two $\gamma$-insertions $w,w_{j}$ merge (following item 1.), and this is integrable. So it suffices to remove a neighbourhood of the diagonal to get the result for $\gamma\;\geqslant\;\sqrt{2}$.
@@
By Proposition 2.6, we have in $\mathcal{C}^{\prime}$, for $\alpha<\alpha_{2,1}$ (recall $\alpha_{-2,1}=\alpha_{2,1}+2\gamma$):
$\mathcal{I}_{2,(1,1),(1,1)}(\alpha)$ $=\Psi_{\alpha_{-2,1}}\int_{\mathbb{D}^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$
$\quad+\int_{\mathbb{D}^{2}}\left(\Psi_{\alpha}(w_{1},w_{2})-|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\right)\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$
$=\Psi_{\alpha_{-2,1}}J_{1}(\alpha)+\int_{\mathbb{D}^{2}}O\left(|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}(|w_{1}|\vee|w_{2}|)^{\xi}\right)\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}},$
for some $\xi>0$, and the notation $J_{1}(\alpha)$ is from Section 2.4.1. The second integral is absolutely convergent and analytic in the neighbourhood of $\alpha_{2,1}$. Thus,

### v2

@@ 2.4.2. Fusion estimates
We now turn to the fusion estimates, which are key to the factorisation of the residues as a product of a primary field and a Dotsenko–Fateev integral. These type of estimates are not new, see e.g. [KRV19, Section 5.1].
The next proposition gives some fusion estimates when we send a number $r\in\mathbb{N}^{*}$ of $\gamma$-insertions to zero. We will only be needing the result for $r\;\leqslant\;3$, but since the proof is similar in all cases, we prefer to state it for arbitrary $r$. Recall that $\mathcal{C}^{\prime}$ is the continuous dual of $\mathcal{C}$.
@@
Let $r\in\mathbb{N}^{*}$, and $\alpha<Q$. The following estimates hold in $\mathcal{C}^{\prime}$.
@@
    Suppose $\alpha+r\gamma\;\geqslant\;Q$. Uniformly over the set of non-coinciding points $\boldsymbol{w}=(w_{1},...,w_{r})\in(\mathbb{D}\setminus\{0\})^{r}$:
@@
    Let $\Delta^{r}:=\{\boldsymbol{w}\in(\mathbb{D}\setminus\{0\})^{r}|\,\forall j=1,...,r-1,\,|w_{j}|\;\geqslant\;e|w_{j+1}|\}$. There exists $\xi>0$ such that, uniformly in the region $\Delta^{r}$:
    $\Psi_{\alpha}(\boldsymbol{w})=\prod_{j=1}^{r}|w_{j}|^{-\gamma\alpha}\prod_{1\;\leqslant\;k<l\;\leqslant\;r}|w_{k}-w_{l}|^{-\gamma^{2}}\left(\Psi_{\alpha+r\gamma}+O(\max_{j}|w_{j}|^{\xi})\right).$
Remark 1.
In the first online version of this work, the second item was stated without the condition $\boldsymbol{w}\in\Delta^{r}$. This is necessary as soon as $2\gamma\;\geqslant\;Q$ because in this case the estimates break down in the region $|w_{1}-w_{2}|\ll\min\{|w_{1}|,|w_{2}|\}$. We thank a careful referee for spotting the mistake.
@@
By permutation symmetry, we can assume $|w_{1}|\;\geqslant\;|w_{2}|...\;\geqslant\;|w_{r}|$ (which is already implied in the second item of the proposition), and we set $t_{j}:=-\log|w_{j}|$ for each j=1,,rj=1,...,r. Let $F\in\mathcal{C}$. By the Girsanov transform, we can write
$\mathbb{E}[\Psi_{\alpha}(\boldsymbol{w})F]$ $=e^{(\alpha+r\gamma-Q)c}\prod_{j=1}^{r}|w_{j}|^{-\gamma\alpha}\prod_{1\;\leqslant\;k<l\;\leqslant\;r}|w_{k}-w_{l}|^{-\gamma^{2}}$
$\qquad\times\mathbb{E}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}}F\left(c+\varphi+\gamma\sum_{j=1}^{r}G(w_{j},\cdot\,)\right)\right]$
Our goal will be to bound the expectation in the last formula. We will first assume that $F=1$ and then explain at the end what are the changes.
Recall the decomposition already used in the proof of Proposition 2.2: we have $X(e^{-t+i\theta})=B_{t}+\varphi_{t}(e^{i\theta})$, where the circle average process $B_{t}=\int_{0}^{2\pi}X(e^{-t+i\theta})\frac{\mathrm{d}\theta}{2\pi}$ is distributed as a standard Brownian motion independent of $(\varphi_{t})_{t\;\geqslant\;0}$. We will write $Y(e^{-t+i\theta}):=\varphi_{t}(e^{i\theta})$. Away from 0, $Y$ differs from $X$ by a continuous function, so its GMC is almost surely well-defined in each annulus $\mathbb{A}_{t}=\{e^{-t}<|z|<1\}$, and we define $Z_{t}:=\lim_{\epsilon\to 0}\int_{\mathbb{A}_{t}}\epsilon^{\frac{\gamma^{2}}{2}}e^{\gamma Y_{\epsilon}(z)}|\mathrm{d}z|^{2}$. Note that $t\mapsto Z_{t}$ is a.s. continuous and increasing on $\mathbb{R}_{+}$, so $\mathrm{d}Z_{t}$ is a (random) measure on $\mathbb{R}_{+}$. Then, the GMC measure of $X$ satisfies $M_{\gamma}(\mathbb{A}_{t})=\int_{0}^{t}e^{\gamma B_{s}}\mathrm{d}Z_{s}$.
@@
$\mathbb{E}\left[\exp\left(-\mu e^{\gamma c}\int_{\mathbb{D}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}\right)\right]$ $\;\leqslant\;\mathbb{E}\left[\exp\left(-\mu e^{\gamma c}\int_{\mathbb{A}_{t_{1}-1}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}\right)\right]$ (2.11)
$\;\leqslant\;C\mathbb{E}\left[\exp\left(-\mu e^{\gamma c}\int_{t_{1}-2}^{t_{1}-1}e^{\gamma(B_{t}+\alpha+r\gamma-Q)t}\mathrm{d}Z_{t}\right)\right]$
Applying Cameron–Martin’s theorem to the drifted Brownian motion inside the last line, we get that the last expectation in the previous displayed equation equals
$\mathbb{E}\left[e^{(\alpha+r\gamma-Q)B_{t_{1}-2}-\frac{1}{2}(\alpha+r\gamma-Q)^{2}(t_{1}-2)}\exp\left(-\mu e^{\gamma(c+B_{t_{1}-2})}\int_{0}^{1}e^{\gamma(\widetilde{B}_{t}+(\alpha+r\gamma-Q)t)}\mathrm{d}Z_{t+t_{1}-2}\right)\right]$
Now, we disintegrate with respect to the law of $B_{t_{1}-2}$ and use that the Gaussian density is bounded by $(2\pi(\log\frac{1}{|w_{1}|}-2))^{-1/2}$ to get that the previous displayed equation is bounded by a constant times
$\frac{|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}}{|\log|w_{1}||^{1/2}}\mathbb{E}\left[\int_{\mathbb{R}}e^{(\alpha+r\gamma-Q)x}\exp\left(-\mu e^{\gamma(c+x)}\int_{0}^{1}e^{\gamma(\widetilde{B}_{t}+(\alpha+r\gamma-Q)t)}\mathrm{d}Z_{t+t_{1}-2}\right)\mathrm{d}x\right].$
Finally, by the stationarity of the process $(Z_{t})_{t\;\geqslant\;0}$, we can replace $Z_{t+t_{1}-2}$ in the last displayed equation by $Z_{t}$. Moreover, we can evaluate the integral over $x\in\mathbb{R}$ using the formula $y^{-s}=\frac{1}{\Gamma(s)}\int_{\mathbb{R}}e^{-ty}t^{s}\frac{\mathrm{d}t}{t}$ for all $y,s>0$. Plugging these estimates into (2.11), we have just proved that
$\mathbb{E}\left[\exp\left(-\mu e^{\gamma c}\int_{\mathbb{D}}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}}\right)\right]$
$\quad\;\leqslant\;C\frac{|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}}{|\log|w_{1}||^{1/2}}\mathbb{E}\left[\left(-\mu e^{\gamma c}\int_{0}^{1}e^{\gamma(\widetilde{B}_{t}+(\alpha+r\gamma-Q)t)})\mathrm{d}Z_{t}\right)^{-\frac{1}{\gamma}(\alpha+r\gamma-Q)}\right]=O(|w_{1}|^{\frac{1}{2}(\alpha+r\gamma-Q)^{2}}).$
Finally, for an arbitrary $F\in\mathcal{C}$, we can assume that $F>0$ up to decomposing into positive and negative parts. Then, we can fix $N>0$ such that $F$ depends only on the first $N$ Fourier modes of the boundary field. Conditionally on these first modes, $F$ is a constant, so we can apply the previous estimates (including Kahane’s convexity inequality). We can integrate out these last modes at the very end and the estimate remains unchanged.
@@
For non-coinciding points $w_{1},...,w_{r}\in\mathbb{D}^{r}$, we introduce the notation
$I_{D}(\boldsymbol{w}):=\int_{D}\prod_{j=1}^{r}|z-w_{j}|^{-\gamma^{2}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma\alpha}},$
for every open set $D\subset\mathbb{D}$. This integral admits positive moments of order $p>0$ for all $p<\min\{\frac{4}{\gamma^{2}},\frac{2}{\gamma}(Q-\alpha-r\gamma)\}$ [KRV20, (2.14)]. We will assume $p<1$ so that the condition is just $p<\frac{2}{\gamma}(Q-\alpha-r\gamma)$, since $\frac{4}{\gamma^{2}}>1$. By an elementary computation of the covariance, we have $X|_{e^{-t}\mathbb{D}}(e^{-t}\,\cdot)\overset{\text{law}}{=}X+\sqrt{t}\delta$, with $\delta$ an independent standard Gaussian. From this observation and the change of variable $z\mapsto e^{t}z$ in the definition of $I_{e^{-t}\mathbb{D}}$, we get the scaling relation $\mathbb{E}[I_{e^{-t}\mathbb{D}}(\boldsymbol{w})^{p}]=e^{-t\xi(p)}\mathbb{E}[I_{\mathbb{D}}(e^{t}\boldsymbol{w})^{p}]$, with $\xi(p)=\gamma(Q-\alpha-r\gamma)p-\frac{\gamma^{2}}{2}p^{2}$. The number $\xi(p)$ is called a multifractal exponent (see [RV14, Theorem 2.14] for more on this topic). We note that $\sup_{\boldsymbol{w}\in\Delta^{r}}\mathbb{E}[I_{\mathbb{D}}(\boldsymbol{w})^{p}]<\infty$ since the $\gamma$-insertions are far away from each other on $\Delta^{r}$, compared to their distance to 0.
@@
Our goal is to control $\mathbb{E}[e^{-\mu I_{\mathbb{D}}(\boldsymbol{w})}-e^{-\mu I_{\mathbb{D}}(0)}]$. First, we show that we can remove the contribution of the disc of radius $|w_{1}|^{1-\eta}$ around 0, for any $\eta\in(0,1)$. Using the elementary inequality $1-e^{-x}\;\leqslant\;x^{p}$ for all $x>0$ and $p\in(0,1)$, we have
$0\;\leqslant\;\mathbb{E}\left[\exp\left(-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})\right)-\exp\left(-\mu I_{\mathbb{D}}(\boldsymbol{w})\right)\right]$ $=\mathbb{E}\left[\exp\left(-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})\right)\left(1-\exp\left(-\mu I_{|w_{1}|^{1-\eta}\mathbb{D}}\right)\right)\right]$
$\;\leqslant\;\mathbb{E}\left[1-\exp\left(-\mu I_{|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})\right)\right]$
$\;\leqslant\;\mathbb{E}[(I_{|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w}))^{p}]=O(|w_{1}|^{(1-\eta)\xi(p)}),$
uniformly over $\Delta^{r}$. Similarly, we get the bound $\mathbb{E}[e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)}-e^{-\mu I_{\mathbb{D}}(0)}]=O(|w_{1}|^{(1-\eta)\xi(p)})$.
It remains to bound the contribution of the annulus $\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}$. In this annulus, we have the estimate $\prod_{j=1}^{r}\frac{|z-w_{j}|^{-\gamma^{2}}}{|z|^{-\gamma^{2}}}=1+O(|w_{1}|^{\eta})$ uniformly as $w_{1}\to 0$. In other words, the two measures in the definitions of $I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})$ and $I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)$ are mutually absolutely continuous, with Radon–Nikodym derivative of the form $1+O(|w_{1}|^{\eta})$ uniformly. Equivalently, there exists $C>0$ such that
$|I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})-I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)|\;\leqslant\;C|w_{1}|^{\eta}I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)\;\leqslant\;C|w_{1}|^{\eta}I_{\mathbb{D}}(0).$
It follows from the inequality $|e^{-x}-e^{-y}|\;\leqslant\;|x-y|^{p}$ valid for $x,y\;\geqslant\;0$ and $p\in(0,1)$, that
$|\mathbb{E}[e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(\boldsymbol{w})}-e^{-\mu I_{\mathbb{D}\setminus|w_{1}|^{1-\eta}\mathbb{D}}(0)}]|\;\leqslant\;C|w_{1}|^{\eta p}\mathbb{E}[I_{\mathbb{D}}(0)^{p}].$
For the choice of $p$ above, the moment is finite, and we are done.
@@
In this proof, we use the following subsets of $\mathbb{D}^{2}$:
$\mathcal{D}_{0}:=\{|w_{2}|<|w_{1}|<e|w_{2}|\}\qquad\text{and}\qquad\mathcal{D}_{1}:=\{|w_{1}|\;\geqslant\;e|w_{2}|\}.$
Note that $\mathcal{D}_{1}=\Delta^{2}$ from Proposition 2.6, and recall the notation $\mathbb{A}_{t}=\{e^{-t}<|z|<1\}$. We will need to treat the cases $\gamma<\sqrt{2}$ and $\gamma>\sqrt{2}$ separately.
Case $\gamma>\sqrt{2}$. In this case, $\alpha_{-2,1}=\alpha_{2,1}+2\gamma>Q$.
as $r\to 0$. At $\alpha=\alpha_{2,1}$, the exponent is $\frac{1}{2\gamma^{2}}(\gamma^{2}-2)^{2}>0$. By continuity in $\alpha$, this exponent is positive in a neighbourhood of $\alpha_{2,1}$. Using as before the covering $\mathcal{D}_{0}\subset\cup_{n\in\mathbb{N}}e^{-n}\mathbb{A}_{2}^{2}$, we see that $\int_{\mathcal{D}_{0}}\Psi_{\alpha}(w_{1},w_{2})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$ is bounded by an absolutely convergent series.
for some $\xi>0$. This is integrable in a neighbourhood of $\alpha_{2,1}$, so we are done.
@@
By Proposition 2.6 and permutation symmetry of $(w_{1},w_{2})$, we have in $\mathcal{C}^{\prime}$ for $\alpha<\alpha_{2,1}$:
$\mathcal{I}_{2,(1,1),(1,1)}(\alpha)$ $=2\Psi_{\alpha+2\gamma}\int_{\Delta^{2}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$ (2.12)
$\quad+2\int_{\Delta^{2}}\left(\Psi_{\alpha}(w_{1},w_{2})-|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\Psi_{\alpha+2\gamma}\right)\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$
$\quad+2\int_{\mathcal{D}_{0}}\Psi_{\alpha}(w_{1},w_{2})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}.$
By the second item of Proposition 2.6, the second line of the right-hand-side is uniformly integrable (hence analytic) up to $\alpha<\alpha_{2,1}+\xi$ for some $\xi>0$. Moreover, using that $\mathcal{D}_{0}\subset\cup_{n\in\mathbb{N}}e^{-n}\mathbb{A}_{2}^{2}$ and the multifractal scaling relation (from the proof of Proposition 2.6), we see that the third line is bounded by an absolutely convergent series, up to a neighbourhood of $\alpha_{2,1}$. Hence, the third line is regular at $\alpha_{2,1}$. Finally, we can argue similarly that the function $\alpha\mapsto\int_{\mathcal{D}_{0}}|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha}|w_{1}-w_{2}|^{-\gamma^{2}}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$ is bounded by an absolutely convergent series in a neighbourhood of $\alpha_{2,1}$, hence it is regular at $\alpha_{2,1}$. Thus, as far as the residue is concerned, we can replace the domain of integration by $\mathbb{D}^{2}$ (instead of $\Delta^{2}$) in the first line of the right-hand-side of (2.12). Hence, by definition of $J_{1}$ (2.9),